[10] 7. Use the Annihilator method to find a particular solution of the equation y – 4y + 4y = 2e27
Answer #1

given 22 A de - y - 4y + 4y = 2e 6-2²y = 26 = 2e2x - where D = d. this is non homogenou, dinear differential Equation of seco6-9) 202x. 00-0) (0-0)2y 0 (8-8) 34 so for the general__solution of this equation like auxillary (n-3)3 = 0 equetónis me 22,2for any doubt regards question comment here I will help you thanks